OpenStudy (anonymous):

I am in Algebra, and i am working on Permutations. I am soooooo lost on these things one of the examples is How many ways can a club of 10 students pick a president, vice president, and treasurer? Lessons answer says : P(10,3) = 720

OpenStudy (amistre64):

10 *9 *8 = 90*8 = 720

OpenStudy (amistre64):

or is it.... 10! 10.9.8 --- = ------- = 5.3.8 = 108 3!7! 3.2

OpenStudy (amistre64):

well if the answer says 720; i guess i was right lol

OpenStudy (amistre64):

i get lazy about halfway thru and start to make up my own question ;)

OpenStudy (anonymous):

sorry my mom needed something.. ^^ thanks amistre! :)

OpenStudy (anonymous):

There are 10 ways to pick the first person, 9 ways to pick the second, and 8 ways to pick the third. So to do all three, there are 10*9*8=720 different possible outcomes.